Y=3 (x+5) (x-2) which of the following formulas is equivalent to this one? Y=3x2+9x-30, or y=x2+3x-10, or y=3x2+3x-10, or y=3x2+3x-30
Answer by jim_thompson5910(35256)


y = 3(x+5)(x-2)

y = 3((x+5)(x-2))

y = 3(x(x-2)+5(x-2))

y = 3(x^2-2x+5x-10)

y = 3(x^2+3x-10)

y = 3x^2+9x-30